The question is
A vending machine takes only nickels and dimes. there are 5 times as many dimes as nickels in the machine. The face value of the coins is 4.40 How many of each coin were in the machine.
I have an equation set up x for nickels and 5y for dimes = 440

Answer by mananth(16949)

A vending machine takes only nickels and dimes. there are 5 times as many dimes as nickels in the machine. The face value of the coins is 4.40 How many of each coin were in the machine.
nickels --------- x
dimes ----------- 5x
..
Value = 440 cents
5x+10*5x=440
5x+50x=440
55x=440
/55
x= 440/55
x= 8 nickesl
dimes = 5x = 40
..
CHECK
40*10 +8*5= 440
...
m.ananth@hotmail.ca